OpenStudy (anonymous):

Convert 1111two to base ten

OpenStudy (anonymous):

What base is it in now?

OpenStudy (campbell_st):

this is \[1 \times 2^3 + 1 \times 2^2 + 1 \times 2^1 + 1 \times 2^0 =\] just evaluate

OpenStudy (anonymous):

so is 15 the answer or is it the way you wrote it out

OpenStudy (anonymous):

i don't really understand

OpenStudy (campbell_st):

ok... out number system is